OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4.

OpenStudy (smarty):

-3y=-3x+4 Divide both sides by -3 y=x+4/-3 Slope= 1 Y Int= -4/3

OpenStudy (ujjwal):

The slope of a line Ax+By+c=0 is -A/B And to find the y intercept put x=0 , in the given equation. Then the value you get for y will be y intercept.